please help me solve this:
find the nth term of the following harmonic series
1/7, 1/11, 1/15, ... n=6

Answer by stanbon(75887)

please help me solve this:
find the nth term of the following harmonic series
1/7, 1/11, 1/15, ... n=6
-----
Only the denominators are changing.
----
Form of the denominators:: 3+4n where n is 1,2,3,...
6th term is 1/(3+4*6) = 1/27
-----
nth term is 1/(3+4n)
--------------------
Cheers,
Stan H.
--------------